After an airplane takes off, it travels 1 km north, and then 2 km east. How far and in what direction is it from the take-off point?



1
Expert's answer
2022-03-29T16:46:17-0400

The magnitude of displacement

"d=\sqrt{1^2+2^2}=2.24\:\rm km"

The direction

"\theta=\tan^{-1}\frac{1}{2}=27^\circ\quad \rm(N\; of\; E)"